# Unveiling Latent Causal Rules: A Temporal Point Process Approach for Abnormal Event Explanation

### 03 Feb 2024 v1

@@ 5.1 E-step: Compute Posterior $Q^{new}(\bm{z}_{i})$
To derive the lower bound of the log-likelihood, let’s denote the a distribution for $\bm{z}_{i}=[z_{ih}]_{h=1,\dots,H}$ as $Q(\bm{z}_{i})$ and $Q\left(\bm{z}_{i}\right)\geq 0$ and $\sum_{\bm{z}_{i}}Q\left(\bm{z}_{i}\right)=1$. The above constraints make $Q(\bm{z}_{i})$ a valid probability mass function.
Now we can create a lower bound for log-likelihood function
@@ 5.2 M-step: Optimize $\theta^{\text{new}}$
For the M-step, we aim to update the model parameters $\theta$ (including the rule set $\mathcal{F}$) to maximize the expected log-likelihood computed in the E-step. which is a lower bound of the log-likelihood, i.e.,
$\theta^{new}:=\arg\max_{\theta}\underbrace{\mathbb{E}_{Q^{new}(\bm{z}_{i})}\left[\log\frac{p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)}{Q^{new}\left(\bm{z}_{i}\right)}\right]}_{\text{``energy"}}$
@@
where $\log p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)$ is the complete data log-likelihood. We further partition $\theta$ as $\theta:=[\theta_{0},\mathcal{F}]$ where $\theta_{0}$ contains the model parameters excluding $\mathcal{F}$. Then we rewrite the above formulation as
@@
5.2.1 Optimization wrt $\mathcal{F}$
We first discuss how to update $\mathcal{F}$ given current $\theta_{0}.$ We aim to propose a differentiable method to learn the rule set. To achieve this goal, we will employ three-fold approximations or reparameterization:
@@ 5.2.1 Optimization wrt $\mathcal{F}$
    Continuous Relaxation of the Rule Learning: We formulate the rule learning problem as a logic variable selection problem and encode the selection results into a binary matrix $A\in\{0,1\}^{H\times(|\mathcal{X}|+M)}$. We then perform the continuous relaxations and relax $A$ to the selection probability $\tilde{A}\in[0,1]^{H\times(|\mathcal{X}|+M)}$.
@@
    Differentiable Top-K Subset Sampling: We perform a differentiable top-K subset sampling with Gumbel noise injected to learn each rule. The learnable parameters are now becoming an unbounded weight matrix $W\in(R^{+})^{H\times(|\mathcal{X}|+M)}$. We leverage simulated annealing by tuning the temperature, which balances the approximation errors and issues caused by gradient vanishing in training.
In summary, we have performed the following series of approximation:
@@
and in the end we optimize $W$ end-to-end in a differentiable way. We provide a more detailed explanation as follows.
I: Continuous Relaxation of the Rule Learning
We first formulate the rule learning as a variable selection problem and introduce a binary matrix $A$ to encode the rule content (with columns indicating the logic variables and rows indicating the rules):
@@
To accommodate for various rule lengths and meanwhile ensure that the rule length dones not exceed $K$, we append $M$ (a hyperparameter) dummy columns after $X_{|\mathcal{X}|}$, indicating empty-predicates to adjust for various rule lengths. We let $\sum_{j=1}^{|\mathcal{X}|+M}a_{hj}=K$, and whenever $a_{hj}=1$ for $j>|\mathcal{X}|$, it means that the rule length, determined by counting the valid predicates, will be less than $K$. A better illustration can be found in Fig. 2. Learning such $A$ involves challenging discrete optimization. Instead, we will relax the problem to learning the selection probability matrix $\tilde{A}$ with $0\leq\tilde{a}_{hj}\leq 1$.
Figure 2: Illustration of rule content. Top diagrams indicate $H$ learned rules given the bottom binary matrix $A$.
Given the selected predicates encoded in $A$, we can further identify their temporal relations. For each paired $X_{u}$ and $X_{v}$, we introduce a learnable parameter $\bm{\alpha}$ to denote their corresponding probabilities
@@
where $\alpha_{b}+\alpha_{e}+\alpha_{a}+\alpha_{n}=1$ and $0\leq\alpha_{b},\alpha_{e},\alpha_{a},\alpha_{n}\leq 1$. That is, $\bm{\alpha}$ belongs to a 3-d simplex. We soften the relation $r_{i}(X_{u},X_{v})$ by replacing it with
@@
For each rule, we consider the temporal relations between all possible pairs of selected predicates. If a rule involves multiple combinations of two distinct predicates, we employ the softmin operation on these $r_{i}$ values. Using softmin helps prevent excessive value reduction resulting from multiple multiplications. The softmin value for a rule $h$ is denoted as ${{\mathcal{T}}}_{h},h=1,\dots,H$. The combination of softmax and softmin operations ensures that the probability learning process remains differentiable. We represent the selection probabilities for all temporal relations in a rule as $\bm{\alpha_{h}}$, and these learned values determine the types of temporal relations.
II: Continuous Approximation of the Boolean Logic Features
To learn the rule set in a differentiable way, in the M-step we propose an approximation to the boolean logic feature as shown in Eq. (3) and rewrite the intensity function as
$\lambda_{\text{soft}}^{*}(t_{i}\mid\bm{z}_{i})=b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}{{\mathcal{T}}}_{h}\mathcal{K}_{Lap}\left\{\sum_{j=1}^{|\mathcal{X}|}X_{j}(t_{j})\tilde{a}_{hj},K\right\}.$ (13)
We use Laplace kernel denoted as $\mathcal{K}_{Lap}$ with center $K$ to reparameterize the boolean feature and make it differentiable with $\tilde{A}$. As previously mentioned, ${{\mathcal{T}}}_{h}$ provides a softened approximation for temporal relations. This approach transforms the complete data log-likelihood into a deterministic function of $\tilde{A}$, enabling the gradient computation through automatic differentiation.
@@
Given the above description, now the rule learning process can be regarded as a top-K subset sampling process, which is governed by the selection probability $\tilde{A}$. We further use the reparameterization trick [18] for this top-K subset sampling.
Assume the top-K subset sampling can be parameterized by an unbounded weight matrix $W=[w_{1};w_{2};\dots;w_{H}]\in(R^{+})^{H\times(|\mathcal{X}|+M)}$, with each element non-negative. The top-$K$ relaxed subset sampling (using Gumbel noise) is summarized in Algorithm 1 (Appendix Section 1). By injecting the Gumbel noise, the introduced key $\hat{r}_{j}$ is differentiable with respect to the input weight $w_{h}$. We can use a top-$K$ relaxation based on successive applications of the softmax function [13] to obtain the relaxed top-K vector $\tilde{a}_{h}$.
To optimize sampling matrix $W$, we have
@@
The parameter $W$ is updated with gradient descent.
With this reparameterization trick, the level of discreteness of $A$ can be controlled by the temperature parameter $\tau$. Thus, in our algorithm, an almost discrete matrix $A$ is used to select rules, and weight matrix $W$ is updated in every iteration to sample $A$. During the learning process, we employed cycled simulated annealing to fine-tune the temperature parameter $\tau$. The choice of $\tau$ balances the trade-off between approximation errors and gradient vanishing issues. Each time when we iterate all data, $\tau$ anneals from a large value to a small non-zero one.
5.2.2 Optimization wrt $\theta_{0}$
M-step: We optimize other model parameters by taking gradient of $\ell(\theta_{0},W)$ wrt $[b_{0},[r_{h}]_{h\in[1,\dots,H]},[\bm{\alpha_{h}}]_{h\in[1,\dots,H]},\bm{\pi}]$, and set these gradients to 0. We need to make sure that these model parameters are non-negative. This can be guaranteed by projected gradient descent or using the change-of-variable trick (e.g., parameterize $b_{0}$ as $\exp(b_{0})$).
@@ 6.1 Synthetic Data Experiments
We assess our model’s learning accuracy in terms of rule discovery, rule weight estimation, and rule prior distribution learning across various scenarios.
@@
In our empirical analysis, we systematically explore various aspects to evaluate the scalability and performance of our method: (i) Diverse Ground Truth Rules: We vary the number of ground truth rules, ranging from 1 to 4, to examine the model’s performance across different complexity levels. (ii) Variation in Predicate Set: We consider five scenarios with different size of predicate set, including 10, 15, 20, 25, and 30, to assess the scalability of our method. A larger logic variable set represents a more challenging problem. (iii) Variation in Data Size: We conduct experiments with varying data sizes, including sequences of 5000, 10,000, 20,000, 30,000, and 40,000 instances, to investigate the EM algorithm’s scalability under different data scales.
It is important to note that only a small subset of the predicates is excluded in the ground truth rules. In other words, the majority of the candidate logic variables are dummy variables. We aim to demonstrate that our method excels in pinpointing the most critical logic variables during the rule learning process. We present the primary results based on experiments with 20,000 sequences for each scenario. For additional details and comprehensive findings, please refer to the Appendix (Section 2).
Figure 3: Our proposed model’s performance is evaluated across all five scenarios for group-2, with two ground truth rules. We evaluate our model’s performance in terms of rule discovery, rule weight learning, and rule prior distribution learning. The color “blue” indicates ground truth rules, weights, and prior distributions, whereas the colors “red” and “yellow” indicate the learning results.
Figure 4: Jaccard similarity score and MAE of rule weights and rule preferential probabilities for all 4 groups. X-axis indicates the predicate library size and Y-axis indicates the value of Jaccard similarity and MAE.
@@
We evaluate the accuracy of our method in terms of rule set discovery and point process model parameter learning. In particular, we assess its accuracy in increasingly challenging scenarios by gradually enhancing the number of dummy logic variables in the predicate set.
In Fig. 3, we present the results for group-2, while the complete results are available in the Appendix (Section 2). The top row of plots employs Venn diagrams to visually depict the agreement between the actual rule set and the learned rule set, quantified through the Jaccard similarity score (the intersection area divided by the union area). Our proposed model consistently identifies almost all the ground truth rules across all groups.
The middle row diagrams compare the true rule weights with the learned rule weights, and the bottom row diagrams compare the true rule prior distributions with the learned rule distributions. Once the ground truth rules are accurately learned, the learned prior rule distributions closely align with the true values, resulting in low mean absolute errors (MAEs).
In addition, we also compared our method with several classic unsupervised rule mining methods, including Apriori [1], NEclatcloesed [2], CM-SPADE [6] and VGEN [7]. These baseline methods identify rules using the frequency thresholds, resulting in coarse rule discovery. From the experiment results, these approaches lead to the extraction of a vast number of noisy and erroneous rules, while the few accurate rules tend to remain hidden within this extensive rule set. A complete results and discussion can be found in Appendix (Section 4 and 6).
@@
Fig. 4 illustrates the Jaccard similarity scores and MAE for all groups as the number of redundant predicates increases. Across all four groups, we observe a decrease in Jaccard similarity score and a slight increase in MAE with the growing number of predicates. However, it’s important to note that our model maintains stable and reliable when the size of the predicate set is appropriate (i.e., around 20-30) and an adequate event sequences are available.
Furthermore, our approach exhibits promising results in identifying the causal rules for each event, inferred by the posterior of the latent rule index. Tab. 1 presents the accuracy of the posterior probabilities for each rule. Notably, in scenarios where almost all ground truth rules were successfully learned, the posterior accuracy is exceptionally high. However, in more challenging tasks, such as those in Group-3 and Group-4, where a larger number of ground truth rules are involved, the accuracy of the rule index posterior experiences a slight decrease.
In fact, our proposed model didn’t deviate significantly in the case of erroneously uncovered rules. Fig. 5 provides an example of a learned wrong rule from Group-3, Case-4. In this instance, the discovered inaccurate rule includes only one additional redundant predicate. Similar patterns are observed in other cases as well.
Figure 5: One example of erroneously uncovered rule for Group-3 (3 ground truth rules) Case-4 (23 to-be-searched predicates). Ground truth rule: $Y\leftarrow x_{4}\land x_{5}\land(x_{4}\,\text{Before}\,x_{5})$. Learned rule: $Y\leftarrow x_{1}\land x_{4}\land x_{5}\land(\ x_{4}\,\text{Before}\ x_{5})$. We see that only one more predicate was wrongly excluded.
@@
1 (10preds) 0.8036 0.7794 0.2818 0.2658
2 (15preds) 0.7962 0.7831 0.4180 0.5155
3 (20preds) 0.7979 0.7717 0.4224 0.4959
4 (25preds) 0.7986 0.6821 0.2802 0.5557
5 (30preds) 0.8039 0.6789 0.2675 0.4960
Table 1: Rule index inference accuracy on synthetic data for all 4 groups.
@@
We compared our model with several state-of-the-art baselines in terms of event predictions, using MAE as the evaluation metric for event time prediction. A complete descriptions of these baseline methods can be found in Appendix Section 4. As shown in Tab. 2, our model outperforms all baselines, particularly in complex scenarios like case-5 with 30 predicates across all groups. In other cases, the MAE varies depending on rule learning outcomes, but it consistently matches or surpasses the baselines.
@@
Table 2: Event time prediction MAE on synthetic data for case-5 (30 preds) of all 4 groups.

### 19 Mar 2024 v2

@@ 5.1 E-step: Compute Posterior $Q^{new}(\bm{z}_{i})$
To derive the lower bound of the log-likelihood, let’s denote the distribution for $\bm{z}_{i}=[z_{ih}]_{h=0,1,\dots,H}$ as $Q(\bm{z}_{i})$ and $Q\left(\bm{z}_{i}\right)\geq 0$ and $\sum_{\bm{z}_{i}}Q\left(\bm{z}_{i}\right)=1$. The above constraints make $Q(\bm{z}_{i})$ a valid probability mass function.
Now we can create a lower bound for the log-likelihood function
@@ 5.2 M-step: Optimize $\theta^{\text{new}}$
For the M-step, we aim to update the model parameters $\theta$ (including the rule set $\mathcal{F}$) to maximize the expected log-likelihood computed in the E-step, which is a lower bound of the log-likelihood, i.e.,
$\theta^{new}:=\arg\max_{\theta}\underbrace{\mathbb{E}_{Q^{new}(\bm{z}_{i})}\left[\log\frac{p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)}{Q^{new}\left(\bm{z}_{i}\right)}\right]}_{\text{``energy''}}$
@@
where $\log p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)$ is the complete data log-likelihood. We further partition $\theta$ as $\theta:=[\theta_{0},\mathcal{F}]$ where $\theta_{0}$ contains the continuous model parameters excluding $\mathcal{F}$. Then we rewrite the above formulation as
@@
5.3 Optimization w.r.t. $\mathcal{F}$
We first discuss how to update $\mathcal{F}$ given current $\theta_{0}.$ We aim to propose a differentiable method to learn the rule set. To achieve this goal, we will employ three-fold approximations and reparameterization:
@@
    Continuous Relaxation for Rule Learning: We first formulate the rule learning problem as a logic variable selection problem and encode the selection results into a binary matrix $A\in\{0,1\}^{H\times(|\mathcal{X}|+M)}$. We then perform the continuous relaxations and relax $A$ to the selection probability $\tilde{A}\in[0,1]^{H\times(|\mathcal{X}|+M)}$.
@@
    Differentiable Top-$K$ Subset Sampling: We conduct a differentiable top-$K$ subset sampling by injecting Gumbel noise into $\tilde{A}$ to generate each rule. This approach introduces noise during rule generation, promoting exploration and aiding in avoiding suboptimal solutions. To control the generated rule length, we further choose to use a Top-$K$ subset sampling [31] by ensuring each rule length does not exceed $K$. In this way, the learnable parameters to determine $\mathcal{F}$ are now becoming an unbounded weight matrix $W\in(R^{+})^{H\times(|\mathcal{X}|+M)}$, which can be learned by gradient-descent. Sepcifically, we leverage simulated annealing by tuning the temperature, which balances the approximation errors and issues caused by gradient vanishing in training.
In summary, we have performed the following series of approximations:
@@
and in the end we optimize $W$ end-to-end in a differentiable way to determine $\mathcal{F}$. We provide a more detailed explanation as follows.
I: Continuous Relaxation for Rule Learning
We first formulate the rule learning as a variable selection problem and introduce a binary matrix $A=[a_{hj}]$ to encode the rule content (with columns indicating the logic variables and rows indicating the rules):
@@
To accommodate for various rule lengths and meanwhile ensure that the rule length does not exceed $K$, we append $M$ (a hyperparameter) dummy columns after $X_{|\mathcal{X}|}$, indicating empty-predicates to adjust for various rule lengths. We let $\sum_{j=1}^{|\mathcal{X}|+M}a_{hj}=K$, and whenever $a_{hj}=1$ for $j>|\mathcal{X}|$, it means that the rule length, determined by counting the valid predicates, will be less than $K$. A better illustration can be found in Fig.2.
Learning such a binary $A$ with the constraints, $\sum_{j=1}^{|\mathcal{X}|+M}a_{hj}=K$, $\forall h$, involves challenging discrete optimization. Instead, we will relax the problem to learning the selection probability matrix $\tilde{A}=[\tilde{a}_{hj}]$ with $0\leq\tilde{a}_{hj}\leq 1$.
Figure 2: Illustration of how to encode the rule content into a binary matrix $A$. In the top diagrams, the content for a total of $H$ rules is depicted. In the bottom diagram, the corresponding encoded binary matrix $A$ is demonstrated. The colorful areas in the matrix indicate a value of 1, while the grey areas indicate a value of 0. Each row in the matrix represents a rule, and each column corresponds to a predicate. $\xi_{1}$ and $\xi_{2}$ are dummy predicates, and $K$=3 in this example.
Similarly, we want to encode the temporal relations with the same idea. Given the selected predicates, we can further identify their temporal relations. For each selected paired $X_{u}$ and $X_{v}$, we introduce a learnable parameter $\bm{\alpha}$ to denote their corresponding probabilities
@@
where $\alpha_{b}+\alpha_{e}+\alpha_{a}+\alpha_{n}=1$ and $0\leq\alpha_{b},\alpha_{e},\alpha_{a},\alpha_{n}\leq 1$. That is, $\bm{\alpha}$ belongs to a 3-$d$ simplex, indicating the probability that the temporal relation belongs to “Before”, “Equal”, “After”, or “None”. We represent the selection probabilities for all temporal relations in a rule as $\bm{\alpha_{h}}$, and these learned values determine the types of temporal relations.
II: Continuous Approximation of the Boolean Logic Features
To learn the rule set in a differentiable way in the M-step, we propose to approximate the log-likelihood function as a deterministic function of $\tilde{A}$ and $[\bm{\alpha_{h}}]_{h}$. This is realized by approximating the boolean logic feature (as shown in Eq.(3)) and rewriting the intensity function as
$\lambda_{\text{soft}}^{*}(t_{i}\mid\bm{z}_{i})=b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}\mathcal{K}_{h}{{\mathcal{T}}}_{h}$ (12)
where $\mathcal{K}_{h}$ and ${{\mathcal{T}}}_{h}$ are deterministic functions of $\tilde{A}$ and $[\bm{\alpha_{h}}]_{h}$, respectively.
For example, we choose to compute $\mathcal{K}_{h}$ by Laplace kernel denoted as $\mathcal{K}_{Lap}$, such as
$\mathcal{K}_{h}=\mathcal{K}_{Lap}\left\{\sum_{j=1}^{|\mathcal{X}|+M}X_{j}(t_{j})\tilde{a}_{hj},K\right\}$ (13)
where we use $\mathcal{K}_{Lap}$ with center $K$ to reparameterize the boolean feature and make it differentiable with $\tilde{A}$. The value of this Laplace kernel approaches 1 if $\sum_{j=1}^{|\mathcal{X}|+M}X_{j}(t_{j})\tilde{a}_{hj}=K$, which indicates each selected predicate should be grounded as true (for dummy predicates $\xi$, we just let each $\xi=1$). In this way, we use $\mathcal{K}_{Lap}$ to approximate the boolean feature for the static predicate parts.
Furthermore, we introduce ${{\mathcal{T}}}_{h}$ to provide a softened approximation for temporal relations. We soften the relation $R_{j}(t_{u},t_{v})$ by replacing it with
@@
For each rule, we consider the temporal relations between all possible pairs of selected predicates. If a rule involves multiple combinations of two distinct predicates, we employ the softmin operation on these $R_{j}$ values. Using softmin helps prevent excessive value reduction resulting from multiple multiplications. The softmin value for a rule $h$ is denoted as ${{\mathcal{T}}}_{h},h=1,\dots,H$. The combination of softmax and softmin operations ensures that the probability learning process for $[\bm{\alpha_{h}}]_{h}$ remains differentiable.
Up to this stage, one may also choose to directly optimize $\tilde{A}$ and $[\bm{\alpha_{h}}]_{h}$ by maximizing the expected likelihood, which will be formulated as a constrained optimization problem. However, sampling the rules from the selection probability will encourage exploration and may help get rid of bad suboptimal solutions.
@@
Given the above description, we propose to generate rules given $\tilde{A}$ in the forward pass computation, which is formulated as a top-$K$ subset sampling process. We will use a reparameterization trick for this subset sampling.
Assume that the top-$K$ subset sampling can be parameterized by an unbounded weight matrix $W=[w_{1};w_{2};\dots;w_{H}]\in(R^{+})^{H\times(|\mathcal{X}|+M)}$, with each element non-negative [31]. The top-$K$ relaxed subset sampling (using Gumbel noise) is summarized in Alg.1. By injecting the Gumbel noise, the introduced key $\hat{r}_{j}$ is differentiable with respect to the input weight $w_{h}$. We can use a top-$K$ relaxation based on successive applications of the softmax function [25] to obtain the relaxed top-$K$ vector $\tilde{a}_{h}$, with the summation constraints automatically satisfied. Specifically, temperature $\tau>0$ and it returns a relaxed $K$-hot vector $\tilde{a}_{h}$, such that as $\tau\rightarrow 0$, RelaxedTopK $(\hat{\mathbf{r}},K,\tau)\rightarrow\sum_{j=1}^{K}\operatorname{TopK}(\hat{\mathbf{r}},K)[j]$.
0:  Weight vector $w_{h}=[w_{h1},w_{h2},\dots,w_{hn}]$, subset size $K$, temperature $\tau>0$
0:  Relaxed $K$-hot vector $\tilde{a}_{h}=[\tilde{a}_{h1},\dots,\tilde{a}_{hn}]$, where $\sum_{j=1}^{n}\tilde{a}_{hj}=K,0\leq\tilde{a}_{hj}\leq 1$
  $\mathbf{\hat{r}}\leftarrow[\quad]$
  for $j\leftarrow 1$ to $n$ do
     $u_{j}\leftarrow\text{Uniform}(0,1)$
     $\hat{r}_{j}\leftarrow-\log(-\log(u_{j}))+\log(w_{hj})$
     $\hat{r}$.append($\hat{r}_{j}$)
  end for
  $\tilde{a}_{h}\leftarrow\text{RelaxedTopK}(\mathbf{\hat{r}},K,\tau)$
Algorithm 1 Relaxed Subset Sampling
Overall, to optimize the sampling matrix $W$, we have
@@
where $\frac{\ell(\theta_{0},W)}{\partial\tilde{a}_{h}}$ is computed based on the approximation as discussed in (II), and the latter two partial gradients are obtained based on the differentiable top-K sampling. The parameter $W$ is updated with stochastic gradient descent.
With this reparameterization trick, the level of discreteness of $A$ can be controlled by the temperature parameter $\tau$. Thus, in our algorithm, an almost discrete matrix $A$ is used to select rules, and the weight matrix $W$ is updated in every iteration to sample $A$. During the learning process, we employed a cyclical annealing scheme to fine-tune the temperature parameter $\tau$. The choice of $\tau$ balances the trade-off between approximation errors and gradient vanishing issues. Each time when we iterate all data, $\tau$ anneals from a large value to a small non-zero one.
Rule content parameter $W$ is optimized first until convergences by ignoring the temporal relations. After $W$ converges, we will proceed to optimize $[\bm{\alpha_{h}}]_{h}$ until convergence. The number of rules $H$ and rule length range $K$ can be tuned via cross-validation or predefined with domain knowledge.
5.3.1 Optimization w.r.t. $\theta_{0}$
M-step: We optimize other model parameters by taking gradient of $\ell(\theta_{0},W)$ w.r.t. $[b_{0},[r_{h}]_{h\in[1,\dots,H]},\bm{\pi}]$, and set these gradients to 0. We need to make sure that these model parameters are non-negative. This can be guaranteed by projected gradient descent or using the change-of-variable trick (e.g., parameterize $b_{0}$ as $\exp(b_{0})$).
@@ 6.1 Synthetic Data Experiments
We assess our model’s learning accuracy in terms of rule discovery, rule weight estimation, and rule prior distribution learning across various scenarios. We also evaluate the inference accuracy by identifying the most likely causal rules for each event.
@@
In our empirical analysis, we systematically explore various aspects to evaluate the scalability and performance of our method: (i) Diverse Ground Truth Rules: We vary the number of ground truth rules, ranging from 1 to 4, to examine the model’s performance across different complexity levels. (ii) Variation in Predicate Set: We consider five scenarios with different sizes of predicate sets, including 10, 15, 20, 25, and 30, to assess the scalability of our method. A larger logic variable set represents a more challenging problem. (iii) Variation in Data Size: We conduct experiments with varying data sizes, including sequences of 5000, 10,000, 20,000, 30,000, and 40,000 instances, to investigate the EM algorithm’s scalability under different data scales.
It is important to note that only a small subset of the predicates is included in the ground truth rules. In other words, the majority of candidate logic variables are irrelevant. We aim to demonstrate that our method excels in pinpointing the most critical logic variables during the rule-learning process. We present the primary results based on experiments with 20,000 sequences for each scenario. For details and comprehensive findings, please refer to Appendix Sec.A.
Figure 3: Our proposed model’s performance is evaluated across all five scenarios for group 2, with two ground truth rules. We evaluate our model’s performance in terms of rule discovery, rule weight learning, and rule prior distribution learning. The color “blue” indicates ground truth rules, weights, and prior distributions, whereas the colors “red” and “yellow” indicate the learning results.
Figure 4: Jaccard similarity score and MAE of rule weights and rule prior probabilities for all 4 groups. The X-axis indicates predicate library size and the Y-axis indicates the value of Jaccard similarity and MAE.
@@
We evaluate the accuracy of our method in terms of rule set discovery and TPP model parameter learning. In particular, we assess its accuracy in increasingly challenging scenarios by gradually enhancing the number of irrelevant logic variables in the predicate set.
In Fig.3, we present the results for group-2, while the complete results are available in the Appendix Sec.A. The top row of the figure employs Venn diagrams to visually depict the agreement between the actual rule set and the learned rule set, quantified through the Jaccard similarity score (the intersection area divided by the union area). A rule is deemed correct when both its included predicates and temporal relations are accurately learned. Our proposed model consistently identifies almost all the ground truth rules across all groups.
The middle row of figures compares the true rule weights with the learned rule weights, and the bottom row of diagrams compares the true rule prior distributions with the learned rule distributions. Once the ground truth rules are accurately learned, the learned prior rule distributions closely align with the true values, resulting in low mean absolute errors (MAEs). The MAE is calculated treating weights and priors as vectors for each case.
In addition, we also compared our model with several classic unsupervised rule mining methods, including Apriori [1], NEclatcloesed [3], CM-SPADE [10] and VGEN [11]. These baseline methods identify rules using the frequency thresholds, resulting in coarse rule discovery. From the experiment results, these approaches lead to the extraction of a vast number of noisy and erroneous rules, while the few accurate rules tend to remain hidden within the extensive rule set. A complete results and discussion can be found in Appendix Sec.C and Sec.E.
@@
Fig.4 illustrates the Jaccard similarity scores and MAE for all groups as the number of redundant predicates increases. Across all four groups, we observe a decrease in the Jaccard similarity score and a slight increase in MAE with the growing number of predicates. However, it’s important to note that our model maintains stability and reliability when the size of the predicate set is appropriate (i.e., around 20-30) and adequate event sequences are available.
Furthermore, our approach exhibits promising results in identifying the latent causal rules for each event. For each event, we denote the rule inferred by our learned rule set as $f$ and the true rule as $f^{*}$. Each rule is a binary vector comprised of two parts, one of which tracks rule content, as shown in Eq.11, while the other records temporal relation. For a pair of predicates, we compute the cosine similarity score $\frac{f^{\top}f^{*}}{\left\|f\right\|\left\|f^{*}\right\|}\in\left[0,1\right]$ for each event. We then take the average over all sequences and report the results in Tab.1. Notably, in scenarios where almost all ground truth rules were successfully learned, the similarity score is exceptionally high. However, in more challenging tasks, such as those in Group-3 and Group-4, where a larger number of ground truth rules are involved, the similarity score of the rule index posterior experiences a slight decrease.
We need to address that our discovered rules do not deviate significantly from the true rules even if they don’t match exactly. Fig.5 provides an example of a learned wrong rule from Group-3, Case-4. In this instance, the discovered inaccurate rule includes only one additional redundant predicate. Similar patterns are observed in other cases as well.
Figure 5: One example of inaccurately uncovered rule for Group-3 (3 ground truth rules) Case-4 (23 to-be-searched predicates). Ground truth rule: $Y\leftarrow x_{4}\land x_{5}\land(x_{4}\,\text{Before}\,x_{5})$. Learned rule: $Y\leftarrow x_{1}\land x_{4}\land x_{5}\land(\ x_{4}\,\text{Before}\ x_{5})$. We see that only one more predicate was wrongly excluded.
@@
1 (10preds) 0.8132 0.7890 0.5033 0.4879
2 (15preds) 0.7916 0.7835 0.4903 0.4900
3 (20preds) 0.7992 0.7708 0.4889 0.4738
4 (25preds) 0.7983 0.7011 0.4799 0.4729
5 (30preds) 0.8007 0.6874 0.4329 0.4295
Table 1: Rule index average inference cosine similarity score on synthetic data for all groups and all cases. For each case we repeated the experiment for 20 times to reduce the randomness.
@@
We compared our model with several state-of-the-art baselines in terms of event predictions, using MAE as the evaluation metric for event time prediction. A complete description of these baseline methods can be found in Appendix Sec.C. As shown in Tab.2, our model outperforms all baselines, particularly in complex scenarios like Case-5 with 30 predicates across all groups. In other cases, the MAE varies depending on rule learning outcomes, but it consistently matches or surpasses the baselines.
@@
Table 2: Event time prediction MAE on synthetic data for complex Case-5 (30 predicates) of all 4 groups.